Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
OPERATION-2025
OPERATION-2024
OPERATION-2023
OPERATION-2022
OPERATION-2021
OPERATION-2020
OPERATION-2019
SASE1-2018
SASE1-2017
Experiment data
|
OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn |
Not logged in |
 |
|
|
Message ID: 180
Entry time: 05 May 2023, 07:23
In reply to: 179
|
Author: |
Giuseppe Mercurio |
Group: |
SCS |
Subject: |
Re: Shift summary |
|
|
Actually, it looks like the sync was lost shortly before 23h last night as the jitter went to 0fs and recovered at 1am when a new sync search probably started.
Giuseppe Mercurio, Diego, Martin, Amina wrote: |
achievements:
- inserted the magnet and measured delay scan at the maximum magnetic field with opposite directions
- data analysis showed that runs measured during the afternoon/evening led to sample degradation (the sample shows a dark spot, change in the scattering signal is also cearly visible)
- corrected spatial overlap
- moved to gresh spot and did delay scan at lower OL fluence (Tr = 35% instead of 70%)
problems:
- PP laser phase shifter when optically lock moved from 1018 to 229.8 on its own, and it was realized few hours later. We moved it back to the original value.
|
|
|
|